In my opinion, we ought to stop making our own drums and accept that outside suppliers offer, said Wim Niewindt, managingProduction Needs Financial Advantage (Disadvantage) of Buying the Drums 1. 2. 3. 65,000 drums 90,000 drums 130,000 drums
Answer #1

1.Financial Advantage of Buying Drums from outside supplier = Relevant cost of producing – cost of buying

= Variable Cost of making + supervision cost + rental cost – cost of buying

= (10.70+9.5*70% + 1.50*70%)*65,000 + 58,500 + 175,500 – 21*65,000

= $65,000

i.e. advantage

52.Financial Advantage = 90,000*18.4 + 58,500 + 175,500 – 21*90,000

= $0

i.e. indifferent

3. Financial Advantage = 18.4*130,000 + 58,500 + 175,500 – 21*130,000

= -$104,000

i.e. disadvantage